Today is bobs birthday. In 50 years time bob will be three times as old as he will be in 6 years time.
How old is Bob today.
show your working using algebra

Answer by TimothyLamb(4379)

x + 50 = 3(x + 6)
x + 50 = 3x + 18
2x = 32
x = 16
---
bob is 16
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php